OpenStudy (anonymous):

need help for integration problem ƒdx/1+√x

OpenStudy (anonymous):

Is that the integral of the derivative of 1 + x^1/2 ? ?

OpenStudy (anonymous):

Because the integral and derivative cancel each other out, leaving 1 + x^1/2 What are the limits of the integral?

OpenStudy (anonymous):

1/1 + x^1/2

OpenStudy (anonymous):

integral is undefineid

OpenStudy (anonymous):

\[\int\limits_{?}^{?} 1 + x ^{1/2}\] Is that it?

OpenStudy (anonymous):

Then the answer is jut 1 + x^1/2

OpenStudy (anonymous):

\[\int\limits_{?}^{?}1\div1+\]

OpenStudy (anonymous):

I think he means$$\int\frac1{1+\sqrt x}dx$$

OpenStudy (anonymous):

yes man

OpenStudy (anonymous):

Lol wow thanks oldrin.

OpenStudy (anonymous):

in which case you multiply top and bottom by \(1-\sqrt{x}\) to get:$$\int \frac{1-\sqrt{x}}{1-x}dx=\int\frac1{1-x}dx-\int\frac{\sqrt{x}}{1-x}dx$$

OpenStudy (anonymous):

can ı do this for indefineid integral because

OpenStudy (anonymous):

maybe x=0

OpenStudy (anonymous):

the first integral is trivial:$$\int\frac1{1-x}dx=-\log(1-x)+C$$

OpenStudy (anonymous):

the second integral is easy once we let \(u=\sqrt{x}\):$$2u\ du=dx$$so we get$$\int\frac{u}{1-u^2}\cdot2u\ du=2\int\frac{u^2}{1-u^2}du=2\int\left(\frac1{1-u^2}-1\right)du$$

OpenStudy (anonymous):

by partial fractions we have that$$\frac1{1-u^2}=\frac{A}{1+u}+\frac{B}{1-u}$$where \(A(1-u)+B(1+u)=1\) so we see that \(A+B=1\) and \(-A+B=0\) so that \(A=B=1/2\) and we have:$$\begin{align*}\int\left(\frac1{1-u^2}-1\right)du&=\int\left(\frac12\cdot\frac1{1+u}+\frac12\cdot\frac1{1-u}-1\right)du\\&=\frac12\log(1+u)-\frac12\log(1-u)-u+C\\&=\frac12\log(1+\sqrt{x})-\frac12\log(1-\sqrt{x})-\sqrt{x}+C\end{align*}$$

OpenStudy (anonymous):

so our final result is $$-\log(1-x)+\log(1+\sqrt{x})-\log(1-\sqrt{x})-2\sqrt{x}+C$$

OpenStudy (anonymous):

hmm nah there msut be something wrong there

OpenStudy (anonymous):

I confused a little :D

myininaya (myininaya):

\[\int\limits_{}^{}\frac{1}{1+\sqrt{x}} dx\] Or we could have just started with the sub \[u=\sqrt{x} => u^2=x => 2u du=dx\] \[\int\limits_{}^{}\frac{2u du}{1+u} \] The maybe do another sub like v=1+u so u=v-1 and dv=du \[2 \int\limits_{}^{}\frac{v-1}{v} dv\] You should be able to take it pretty easy for here

OpenStudy (anonymous):

lool that works too

OpenStudy (anonymous):

the error in my approach was a small sign screw up:$$-\log(1-x)-\log(1+\sqrt{x})+\log(1-\sqrt{x})+2\sqrt{x}+C$$

OpenStudy (anonymous):

oh its very easy

OpenStudy (anonymous):

thanks for everyone

OpenStudy (anonymous):

can we use two sub in this question

myininaya (myininaya):

I did.

myininaya (myininaya):

See above.

OpenStudy (anonymous):

yes ı saw thanks a lot

myininaya (myininaya):

You can do multiple subs for any problem

OpenStudy (anonymous):

yes can ı ask you a question

myininaya (myininaya):

You could have wrote the two subs as one sub those

myininaya (myininaya):

\[\int\limits_{}^{}\frac{1}{1+\sqrt{x}} dx\] Let \[v=1+\sqrt{x} =>v-1=\sqrt{x} => (v-1)^2=x => 2(v-1) dv=dx\]

myininaya (myininaya):

\[\int\limits_{}^{}\frac{2(v-1)dv}{v}\]

myininaya (myininaya):

Go ahead, what is the question?

OpenStudy (anonymous):

could we here with the multiplication of the expression x or conjugate for example \[\sqrt{x}-1 \]

myininaya (myininaya):

You mean multiply the conjugate of sqrt(x)+1 on bottom and top? I think that is what the other dude did as a first step.

myininaya (myininaya):

Why do that when you just need a simple sub though

OpenStudy (anonymous):

yes ı mean this is it possible

OpenStudy (anonymous):

the answer is \[2\sqrt{x}+2-2\ln \left| \sqrt{x}+1 \right| + c\] isnt this